# DrishtiKon: Multi-Granular Visual Grounding for Text-Rich Document Images

### v1

@@ Abstract
Visual grounding in text-rich document images is a critical yet underexplored challenge for document intelligence and visual question answering (VQA) systems. We present DhrishtiKon, a multi-granular visual grounding framework designed to enhance interpretability and trust in VQA for complex, multilingual documents. Our approach integrates robust multi-lingual OCR, large language models, and a novel region matching algorithm to accurately localize answer spans at block, line, word, and point levels. We curate a new benchmark from the CircularsVQA test set, providing fine-grained, human-verified annotations across multiple granularities. Extensive experiments demonstrate that our method achieves state-of-the-art grounding accuracy, with line-level granularity offering the best trade-off between precision and recall. Ablation studies further highlight the benefits of multi-block and multi-line reasoning. Comparative evaluations with leading vision-language models reveal the limitations of current VLMs in precise localization, underscoring the effectiveness of our structured, alignment-based approach. Our findings pave the way for more robust and interpretable document understanding systems in real-world, text-centric scenarios. Code and dataset has been made available at https://github.com/kasuba-badri-vishal/DhrishtiKon.
@@ 1 Introduction
Visual Grounding is the task of locating specific visual elements in an image that correspond to a given text query/question in natural language as input to a corresponding probable answer that answers the query from the input image[8]. This task is particularly relevant in the context of Visual Question Answering (VQA) task, where the goal is to provide accurate and contextually relevant answers based on the visual content of an image, providing both transparency and trust on top of the predicted answer’s relavance within the image. Also, Grounded learning for the task allows for more accurate and contextually relevant responses, as it requires models to focus on the visual content directly related to the question to finding the grounded answers. In this work, we will explore Visual Grounding task on text-centric images and also evaluate the performance on a constructed multi-granular grounding benchmark from the collection of Government Circular images. We propose a Multi-granular Visual Grounding framework, DhrishtiKon which aims to enhance the model’s ability to ground answers in visual elements at multiple levels of granularity, supporting multiple pages which utilized a novel region matching algorithm to improve the grounding process.
2 Literature Study
@@
The rise of Multimodal Large Language Models (MLLMs) has advanced document understanding, yet visual grounding in text-rich document images remains an underexplored challenge but yet vital for document intelligence tasks. Recent efforts such as TGDoc[7], DOGE-Bench[9], and TRIG-Bench[3] have introduced benchmarks and models to address this gap. The Table 1 compares these benchmarks across data, methodologies, support and synthetic data generation strategies.
Table 1: Comparison of Visual Grounding Benchmarks on Text-Rich Images
Benchmarks Image Source Quantity Granularity Support Synthetic data generation OCR LLM TGDoc Online Powerpoint samples 99k Slides 12k high quality Chats Single Single-Block PaddleOCR GPT4 DOGE-Bench Charts, Posters, PDFs 700k Instruction 4k Benchmark Multi-Granular Single-Block EasyOCR GPT4o TRIG-Bench DocVQA, InfographicVQA 90k Synthetic 800 Manual Single Multi-Block PaddleOCR GPT4o
@@
TGDoc[7] improves spatial awareness in MLLMs through instruction tuning on 99K PowerPoint slides and 12K GPT-4-generated dialogues, using BLIP2, PaddleOCR, and GPT-4 to align text recognition with spatial cues. DOGE-Bench[9] introduces multi-granular grounding and referring tasks across charts, posters, and PDFs, using a synthetic pipeline (DOGE-Engine) to produce 700K training samples and a 4K evaluation benchmark, enabling fine-grained referencing at block, line, and word levels. In contrast, TRIG-Bench[3] emphasizes grounding in real-world, layout-heavy documents using a hybrid OCR-LLM-human pipeline, offering 90K synthetic and 800 manually curated samples to assess MLLM performance in spatially complex settings. Collectively, these benchmarks expose the limitations of current MLLMs in handling dense layouts and highlight the need for improved spatial reasoning in document-centric tasks. While these benchmarks have made significant strides in visual grounding, they also reveal critical gaps in the current state of MLLMs. The key gaps remain particularly in multilingual, multi-block scenarios with human-verified annotations applicable to all types of text-rich images. In this thesis, we aim to develop multilingual, multi-block and multi-granular visual grounding benchmark that address these limitations and could pave the way for more robust and versatile document understanding systems.
3 DhrishtiKon Methodology
This section details the proposed pipeline for DhrishtiKon, a multi-granular visual grounding solution for text-rich document images. Our approach integrates multi-lingual OCR[4, 5], large language models (LLMs) [2], and a hybrid scoring-based region matching algorithm as discussed in Section 3.4 for accurate localization of semantically relevant text blocks. Figure 1 provides an overview of the DhrishtiKon pipeline.
@@ 3 DhrishtiKon Methodology
Figure 1: DhrishtiKon: Multi-granular Visual Grounding Solution Pipeline with matching algorithm flowchart on an example Illustration VQA sample.
@@ 3.1 Input Acquisition
We begin with an input image of a text-rich document, typically a scanned official circular or memorandum, which forms the visual basis for question answering. Accompanying the image is a natural language question that queries for any extractive information, such as specific data, references, or contextual details present in the document.The question is designed to be open-ended, allowing for a range of possible answers that may span multiple text blocks within the document, but extractive in nature intending to find specific phrases or data points within the image content.
@@ 3.3 Question-Conditioned Answer Prediction
The extracted blocks, along with the user-input question, are fed into a large language model (LLM). We use LlaMA-3.1-8B instruct open-source model[2] as our model, which performs semantic reasoning over the textual content and outputs a predicted answer. This step simulates an MLLM (Multimodal Large Language Model) inference in purely textual modality, bypassing the need for visual input during answer generation as it takes OCR output as contextual information.
@@ 3.4 Candidate Region Matching Algorithm
To locate the predicted answer within the document, we implement a matching algorithm that loops over each probable text block extracted by the OCR engine. For each block, a composite matching score is calculated using the following components:
@@
As described in the Algorithm 1, Each candidate region is scored based on components which inturn have a scale of contribution, and those regions that scored above a pre-defined threshold are retained. The candidate blocks are sorted based on their composite scores. The top-k matched blocks are selected and visualized along with the predicted answer for interpretability. This allows the luxury to choose the number of multi-block predictions required based on the threshold parameters.
Algorithm 1 Region Matching Algorithm for Block-Level Probable Regions
@@ 4.1 Grounding Datasets
Table 2: Statistics of the annotated dataset from CircularsVQA testset for Multi-Granular Visual Grounding.
@@ 4.1.1 Annotation Tool
Figure 2: In-house annotation tool supporting multi-level region labeling.
To construct the annotated test set used in our experiments, we developed a custom in-house annotation tool designed to support hierarchical and fine-grained annotation of document regions. This tool enables annotation at the block, line, word, and point levels, offering flexibility in defining answer spans that match natural visual layouts. The tool has an interactive interface (Figure 2) that allowed to label precise spans and relationships and also supports automatic text detection at word-level from the line-level annotated region aiding in enhanced annotation. Unlike other existing datasets[7, 9], the dataset has been manually annotated instead of using synthetic data generation straties via properietary LLM.
Figure 3: Visual grounding results across different granularities: Block, Line, Word, and Point. The bounding boxes indicate the predicted regions corresponding to the question and answer.
@@ 4.2 Granularity-wise Evaluation
Table 3: Evaluation of grounding accuracy at different granularities
@@ 4.3 Model-wise Evaluation
Input (Text + BBox) OCR (QA) Model (QA) OCR (Grnd.) Model (Grnd.) P (%) R (%) F (%)
@@
Predicted Answer NO PATRAM YES Algorithm 35.58 30.97 33.12
The Region Matching Algorithm, when paired with ground truth answers, clearly outperforms all other methods and could be considered as the skyline performance for line-level grounding tasks in this dataset as the ideal conditions are met.
Qwen2.5-VL[6], despite being a general-purpose multimodal model, underperforms drastically across all metrics. This suggests its current limitations of OCR-free and VLM methods in handling text-centric visual grounding tasks.
@@
We also present a comparative evaluation at the block level in Table 5. Once again, the Region Matching Algorithm maintains superior performance at the block level, emphasizing its consistent advantage in spatially grounded reasoning. The performance gap also reinforces the importance of layout-aware alignment in real-world document QA settings.
Table 5: Block-level grounding evaluation across methods
Method Precision (%) Recall (%) F1-score (%)
Region Matching Algorithm 53.89 68.22 60.21
LLaMA 3.1 (Q+A) 41.72 67.47 51.56
LLaMA 3.1 (Question Only) 32.88 54.46 41.01
@@ 4.4 Ablation Study
To understand the impact of supporting multi-block and multi-line regions in the proposed region matching algorithm, we perform a detailed ablation study over the number of allowable blocks and lines during inference that enhances the accuracy.
@@
Table 6: Effect of max blocks on region matching performance.
@@
We further extended the ablation over the number of lines allowed in the answer span. Table 7 and Figure 4 show the variation in precision, recall, and F1-score across iterations. Precision drops steadily as more lines are aggregated, while recall consistently improves, reaching a plateau around 10 lines. The F1-score stabilizes around 69.3%, with the optimal performance achieved at 5 lines.
Table 7: Effect of max lines on region matching performance.
@@
Figure 4: Precision, Recall, and F1-score trends across increasing max line thresholds.
@@
5 Conclusion and Future works
In this work, we presented DhrishtiKon, a multi-granular visual grounding framework designed to enhance the interpretability and trustworthiness of VQA systems on text-centric images. By integrating multi-lingual OCR, LLMs, and a novel region matching algorithm, we demonstrated significant improvements in grounding accuracy across multiple levels of granularity. Our experiments on a curated dataset derived from CircularsVQA testset showed that mid-level granularities (line-level) yield the best performance for document-based question answering tasks. The ablation studies further validated the effectiveness of multi-block and multi-line reasoning in improving grounding precision and recall. Also, the evaluations highlight that while large vision-language models can produce reasonable semantic predictions, they fall short in precise visual localization tasks without structured grounding mechanisms. Our Region Matching approach demonstrates that traditional alignment-based methods—when combined with robust OCR and textual matching—remain highly competitive for document understanding tasks that demand spatial fidelity.

### v2

# DrishtiKon: Visual Grounding at Multiple Granularities in Documents

@@ Abstract
Visual grounding in text-rich document images is a critical yet underexplored challenge for Document Intelligence and Visual Question Answering (VQA) systems. We present DrishtiKon, a multi-granular and multi-block visual grounding framework designed to enhance interpretability and trust in VQA for complex, multilingual documents. Our approach integrates multilingual OCR, large language models, and a novel region matching algorithm to localize answer spans at the block, line, word, and point levels. We introduce the Multi-Granular Visual Grounding (MGVG) benchmark, a curated test set of diverse circular notifications from various sectors, each manually annotated with fine-grained, human-verified labels across multiple granularities. Extensive experiments show that our method achieves state-of-the-art grounding accuracy, with line-level granularity providing the best balance between precision and recall. Ablation studies further highlight the benefits of multi-block and multi-line reasoning. Comparative evaluations reveal that leading vision-language models struggle with precise localization, underscoring the effectiveness of our structured, alignment-based approach. Our findings pave the way for more robust and interpretable document understanding systems in real-world, text-centric scenarios with multi-granular grounding support. Code and dataset are made available for future research.
@@ 1 Introduction
Visual grounding refers to the task of accurately localizing regions in an image that correspond to a given natural language query or question [23]. In the domain of Document Intelligence and Visual Question Answering (VQA), visual grounding plays a pivotal role in linking textual answers to their supporting evidence within visually complex, text-rich document images—such as government circulars, official memos, and administrative forms. This spatial linkage not only facilitates interpretability and trust but also enables downstream applications that rely on the provenance of extracted information.
While visual grounding has shown significant progress in natural image understanding [23, 13, 16], its application to document images remains underexplored. Document layouts are inherently distinct: they are dense, heterogeneous, often multilingual, and exhibit a variety of hierarchical structures. These characteristics introduce unique challenges for grounding models, particularly in scenarios requiring high precision and accountability—such as compliance auditing, information verification, and explainable AI. Manually tracing the source of an answer in such documents is labor-intensive and prone to human error, emphasizing the need for automated, fine-grained visual grounding methods tailored to Document VQA [11, 19]. In contrast to natural images, where objects are typically grounded spatially, document images require grounding over structured textual units. In this context, the spatial provenance of an answer— where it appears on the document—is as critical as the answer content itself. Accurately identifying and localizing supporting regions enhances model transparency, facilitates human verification, and supports regulatory tasks such as information extraction, digital preservation, and legal compliance.
In this paper, we tackle the task of multi-granular visual grounding in Document VQA. Given a document image and a question, our objective is to return both the textual answer and its precise supporting region(s) in the document at varying levels of granularity. To this end, we propose DrishtiKon, a novel framework that synergizes multilingual OCR, large language models (LLMs), and a hybrid region matching strategy to enable accurate answer localization at multiple granularities—block, line, word, and point. To evaluate our approach systematically, we construct a new benchmark, Multi-Granular Visual Grounding (MGVG) enabling robust evaluation of grounding methods and highlights the challenges posed by real-world administrative documents. Our key contributions are as follows:
  • We introduce DrishtiKon, a multi-granular visual grounding framework tailored for document VQA, supporting localization at block, line, word, and point levels.
  • We design a robust hybrid region matching algorithm that integrates fuzzy text matching, length normalization, and semantic penalties to accurately localize multi-line and multi-block answers.
  • We curate and release the MGVG benchmark, a new dataset of 70 government circulars with 509 diverse QnA pairs and human-annotated fine-grained spatial labels.
  • We perform extensive experiments and ablation studies to quantify grounding performance across granularity levels, offering insights into precision-recall trade-offs in Document VQA with grounding.
  • We benchmark leading vision-language models and highlight their limitations in spatial grounding, demonstrating the superiority of our alignment-based framework for structured document understanding.
2 Related Work
@@
While Multimodal Large Language Models (MLLMs) have advanced document understanding [6, 20, 15, 17, 7, 10, 1], visual grounding in text-rich documents remains a crucial yet underexplored challenge for document intelligence. Recent works such as TGDoc [22], DOGE-Bench [24], and TRIG-Bench [9] have introduced benchmarks and models to address this gap. Table 1 compares these benchmarks in terms of methodologies, granularity, and synthetic data generation strategies.
@@
TGDoc [22] enhances spatial awareness in MLLMs through instruction tuning on 99K PowerPoint slides and 12K GPT-4-generated dialogues, leveraging BLIP2 [8], PaddleOCR [2], and GPT-4 [18] to align text recognition with spatial cues. DOGE-Bench [24] introduces multi-granular grounding and referring tasks across charts, posters, and PDFs, using a synthetic pipeline (DOGE-Engine) to generate 700K training samples and a 4K evaluation benchmark, enabling fine-grained referencing at block, line, and word levels. In contrast, TRIG-Bench [9] focuses on grounding in real-world, layout-heavy documents using a hybrid OCR-LLM-human pipeline, offering 90K synthetic and 800 manually curated samples to assess MLLM performance in spatially complex settings. Collectively, these benchmarks reveal the limitations of current MLLMs in handling dense layouts and highlight the need for improved spatial reasoning in document-centric tasks.
Despite these advances, several critical gaps remain. Existing benchmarks often have limited multilingual coverage, lack comprehensive human-verified annotations, and do not fully address multi-block and multi-line answer localization in complex, real-world documents. Moreover, most prior approaches rely heavily on synthetic data or focus on single-granularity tasks, which limits their applicability to the diverse document types encountered in practice.
Our work addresses these limitations by introducing a multi-granular, multilingual, and human-annotated benchmark, and by proposing a novel region matching algorithm that enables robust answer localization across multiple granularities. This positions our approach as a significant step forward in developing interpretable and trustworthy document understanding systems.
Benchmark Granularity Multi-Block OCR LLM TGDoc Single No PaddleOCR GPT-4 DOGE-Bench Multi No EasyOCR GPT-4o TRIG-Bench Single Yes PaddleOCR GPT-4o Ours Multi Yes
Table 1: Comparison of text-rich visual grounding benchmarks.
3 DrishtiKon Methodology
This section details the proposed pipeline for DrishtiKon, a multi-granular visual grounding solution for text-rich document images. Our approach is designed to address the unique challenges of document intelligence by enabling precise, interpretable answer localization across multiple levels of granularity. The framework integrates advanced multi-lingual OCR [12, 14], state-of-the-art large language models (LLMs) [5], and a novel, hybrid scoring-based region matching algorithm. This combination allows for robust handling of complex layouts, multilingual content, and diverse document structures. Figure 1 provides an overview of the DrishtiKon pipeline. Our methodology is designed for extensibility and can be adapted to a variety of document analysis tasks beyond VQA, including information extraction, compliance checking, and digital archiving. The following subsections describe each component of the pipeline in detail.
@@ 3 DrishtiKon Methodology
Figure 1: DrishtiKon: Multi-granular Visual Grounding Pipeline with matching algorithm flowchart on an example Illustration.
@@ 3.1 Input Acquisition
We begin with an input image of a text-rich document, typically a scanned official circular or memorandum, which forms the visual basis for question answering. The image is accompanied by a natural language question that queries for specific, extractable information, such as data, references, or contextual details present in the document. The open-ended question allows for a range of possible answers spanning multiple text blocks within the document and is extractive, seeking specific phrases or data points in the image.
@@ 3.3 Question-Conditioned Answer Prediction
The extracted blocks, along with the user-input question, are fed into a large language model (LLM). We use LlaMA-3.1-8B instruct open-source model [5] as our model, which performs semantic reasoning over the textual content and outputs a predicted answer. This step simulates multimodal large language model (MLLM) inference in a purely textual modality, bypassing the need for visual input during answer generation by using OCR output as contextual information.
@@ 3.4 Candidate Region Matching Algorithm
To locate the predicted answer, we use a matching algorithm that iterates over each text block extracted by the OCR engine. For each block, a composite matching score is calculated using the following components:
@@
As described in the Algorithm 1, Each candidate region is scored based on components which inturn have a scale of contribution, and those regions that scored above a pre-defined threshold are retained. The candidate blocks are sorted based on their composite scores. The top-k matched blocks are selected and visualized along with the predicted answer for interpretability. This allows selection of the number of multi-block predictions based on the threshold parameters.
Algorithm 1 Region matching algorithm for identifying probable block-level regions: Flowchart illustrating the steps involved in matching regions based on visual and textual features.
@@
3.5 Word-level Grounding Procedure
(a) Step 1: Line-level grounding identifies the relevant line(s) containing the answer.
(b) Step 2: Word-level matching highlights the individual words within the line(s) that correspond to the answer tokens.
(c) Step 3: Extraction of the longest contiguous sequence of matched words forms the final word-level grounded region.
Figure 2: Overview of the word grounding process.
Word-level grounding is designed to achieve fine-grained localization of answer spans by identifying the precise words in a document image that correspond to the predicted answer. This process builds upon the results of line-level grounding and consists of the following steps:
  1. 1.
    Line-to-Word Mapping via Word Match: Starting from the lines identified as relevant by the line-level grounding module (see Figure 2(a)), each line is tokenized into individual words. For each predicted answer, a word-level matching operation is performed to locate all occurrences of the answer tokens within the candidate lines. This ensures that only those words directly contributing to the answer are considered for further localization (Figure 2(b)).
  2. 2.
    Extraction of Longest Contiguous Word Indices: To ensure semantic coherence and minimize fragmentation, the algorithm searches for the longest possible contiguous sequence of matched words that collectively form the answer span. This is achieved by identifying the maximal subsequence of word indices within the line(s) that matches the answer phrase, allowing for minor OCR or tokenization discrepancies. The resulting word-level bounding boxes are then aggregated to produce the final grounded region (Figure 2(c)).
This approach enables highly precise answer localization, supporting downstream applications such as information extraction, document redaction, and explainable VQA.
@@ 4.1 Datasets
Table 2: Dataset statistics for Multi-Granular Visual Grounding (MGVG) test set: Distribution of instance counts across different granularity levels (block, line, word, point) within the annotated dataset.
@@ 4.1.1 Annotation Tool
Figure 3: In-house annotation tool enabling multi-level region labeling: Tool interface showing options for defining and annotating regions of interest at varying levels of granularity.
To construct the annotated test set used in our experiments, we developed a custom in-house annotation tool designed to support hierarchical and fine-grained annotation of document regions. This tool enables annotation at the block, line, word, and point levels, offering flexibility in defining answer spans that match natural visual layouts. The tool has an interactive interface (Figure 3) that allowed to label precise spans and relationships and also supports automatic text detection at word-level from the line-level annotated region aiding in enhanced annotation. Unlike other existing datasets [22, 24], the dataset has been manually annotated instead of using synthetic data generation strategies via proprietary LLM.
@@ 4.2 Granularity-wise Evaluation
Figure 4: Visual grounding performance at varying granularities: Predicted bounding boxes show regions identified for question and answer pairs at block, line, word, and point levels.
@@ 4.3 Model-wise Evaluation
Input (Text + BBox) OCR (QA) Model (QA) OCR (Grnd.) Model (Grnd.) P (%) R (%) F1 (%)
@@
Predicted Answer NO PATRAM [3] YES Algorithm 35.58 30.97 33.12
The Region Matching Algorithm, when paired with ground truth answers, outperforms all other methods and represents the skyline performance for line-level grounding tasks on this dataset under ideal conditions.
Qwen2.5-VL [21], despite being a general-purpose multimodal model, underperforms drastically across all metrics. This suggests its current limitations of OCR-free and Vision-Language Model (VLM) methods in handling text-centric visual grounding tasks.
@@
We also present a comparative evaluation at the block level in Table 5. Once again, the Region Matching Algorithm maintains superior performance at the block level, emphasizing its consistent advantage in spatially grounded reasoning. The performance gap also reinforces the importance of layout-aware alignment in real-world document Question Answering (QA) settings.
Method P (%) R (%) F1 (%)
Algorithm 53.89 68.22 60.21
LLaMA with Ground Truth 41.72 67.47 51.56
LLaMA with Predicted Answer 32.88 54.46 41.01
Table 5: Comparison of block-level grounding accuracy across different methods: Results are reported as IoU on the test set.
4.4 Qualitative Comparison: Limitations of LLM-based and Algorithm-based Grounding
Figure 5 presents a qualitative comparison that highlights the complementary limitations of algorithm-based and LLM-based visual grounding approaches in document VQA. For the example question, ”Where has Shri T.P. Singh been transferred to?”, the algorithm-based method (Figure 5(a)) identifies multiple candidate regions by matching answer tokens and leveraging spatial cues. While effective in high-recall detection, it often results in over-selection, including redundant or irrelevant text regions due to its limited capacity for contextual reasoning. Conversely, the LLM-based method (Figure 5(b)), when provided with OCR-extracted text, demonstrates better semantic understanding of the question and can localize the correct answer span. However, it may fail to capture the complete supporting context, particularly when the answer spans multiple lines or regions, and is sensitive to OCR noise or incomplete textual input. This analysis underscores the inherent trade-offs between the two paradigms—algorithmic methods offer precise spatial grounding but lack semantic depth, while LLMs provide contextual understanding but often struggle with spatial alignment. These observations motivate the need for hybrid approaches that integrate structural precision with semantic reasoning for effective visual grounding in complex document understanding scenarios.
(a) Algorithm-based grounding highlights multiple candidate regions but cannot reason about context, leading to over-selection.
(b) LLM-based [5] grounding predicts the answer but may fail to capture the full supporting content, especially for multi-region answers.
Figure 5: Qualitative comparison of individual LLM-based and algorithm-based grounding (a) Algorithm-based method is precise in matching but lacks reasoning. (b) LLM-based approach can understand context but may miss spatial or structural details. This example demonstrates the complementary limitations of both approaches in Document VQA.
@@ 4.5 Ablation Study
To understand the impact of supporting multi-block and multi-line regions, we perform a detailed ablation study on the proposed region matching algorithm, varying the number of allowable blocks and lines during inference to enhance accuracy.
@@
Table 6: Impact of the maximum number of blocks on region matching performance: Evaluation conducted on the test set, showing the relationship between the number of blocks and IoU.
@@
We further extended the ablation over the number of lines allowed in the answer span. Table 7 and Figure LABEL:fig:line_trend show the variation in precision, recall, and F1-score across iterations. Precision drops steadily as more lines are aggregated, while recall consistently improves, reaching a plateau around 10 lines. The F1-score stabilizes around 69.3%, with the optimal performance achieved at 5 lines.
@@
Table 7: Impact of the maximum number of lines on region matching performance: Results show the relationship between the maximum number of lines considered and the IoU score on a testset.
Figure 6: Trends in precision, recall, and F1-score with increasing maximum line thresholds at IoU = 0.5
@@
4.6 Granularity and Multi-Block Trade-offs: Grounding Quality vs. Complexity
The choice of grounding granularity and the handling of multi-block answers have a significant impact on both the quality of localization and the complexity of text similarity evaluation in document VQA.
Granularity of Grounding:
As illustrated in Figure 7, word-level grounding enables highly precise localization of answer spans, resulting in superior grounding quality. However, this increased precision comes at the cost of higher complexity in text similarity computation, as the system must accurately match and align individual words, often in the presence of OCR noise or tokenization discrepancies. In contrast, block-level grounding simplifies the matching process by considering larger text regions, which reduces the complexity of text similarity evaluation but may lead to less precise localization and lower interpretability.
Figure 7: Comparison of grounding quality and text similarity complexity at different granularities. Word-level grounding achieves higher localization quality but increases text similarity complexity, while block-level grounding simplifies matching at the expense of precision.
Multi-Block Grounding:
Document questions may require aggregating information from multiple, spatially separated blocks. As shown in Figure 8, supporting multi-block grounding can improve answer accuracy for complex queries, but it also increases the complexity of both region selection and text similarity computation. When only a single block is considered, the process is simpler but may miss relevant information, reducing accuracy. Allowing multiple blocks to be selected increases the likelihood of capturing the complete answer, but introduces additional challenges in merging, ranking, and evaluating the candidate regions.
Figure 8: Impact of multi-block grounding on accuracy and complexity. Allowing multiple blocks can improve answer accuracy for complex queries, but increases the complexity of region selection and evaluation.
These trade-offs highlight the importance of selecting an appropriate grounding granularity and region aggregation strategy based on the specific requirements of the document VQA task. Our experiments quantitatively validate these observations, demonstrating that line-level grounding often provides the best balance between localization quality and computational complexity for structured documents.
5 Conclusion
In this work, we presented DrishtiKon, a comprehensive multi-granular visual grounding framework aimed at enhancing the interpretability and reliability of VQA systems on text-rich document images. By integrating advanced multilingual OCR, large language models, and a novel region matching algorithm, our framework achieves accurate answer localization across multiple levels of granularity—block, line, word, and point. To support rigorous evaluation, we introduced the Multi-Granular Visual Grounding (MGVG) benchmark. This benchmark provides a diverse set of human-annotated samples, enabling systematic assessment of visual grounding performance at varying granularities. Through extensive experiments and ablations, we show that mid-level granularities, especially line-level grounding, strike the optimal balance between precision and recall for document VQA. Our results further highlight the necessity of supporting multi-line and multi-block reasoning to handle the structural complexity inherent in real-world documents. Our comparative analysis with state-of-the-art vision-language models reveals their limitations in fine-grained spatial localization, underscoring the advantages of structured, alignment-based approaches like DrishtiKon. Together, our framework and benchmark establish a strong foundation for future research in transparent, accountable, and scalable document understanding systems. Future work will focus on extending the framework to support multi-page and mixed-script grounding where the document is not a single page.
Broader Impact.
The proposed framework holds significant potential for real-world applications such as automated compliance verification, digital recordkeeping, and reliable information extraction across government, legal, and enterprise settings. By improving transparency in automated document analysis, DrishtiKon can contribute to more trustworthy AI systems and foster wider adoption of document intelligence technologies in critical decision-making workflows.
Privacy Considerations.
To safeguard sensitive information and ensure ethical data handling, certain text regions in our document images, particularly those containing names or personally identifiable information, have been blurred. While many source documents originate from publicly available government circulars, we adopt a privacy-preserving approach to maintain responsible data sharing and protect individual and institutional confidentiality.